OpenStudy (anonymous):

find 3v-w given v(2,-3) and w(5,-2)

jimthompson5910 (jim_thompson5910):

v and w look like vectors

jimthompson5910 (jim_thompson5910):

so if v = (2,-3), then 3v = (6, -9) notice I just multiplied everything by 3

jimthompson5910 (jim_thompson5910):

do you know how to take it from here?

OpenStudy (anonymous):

no i dont

jimthompson5910 (jim_thompson5910):

from here, you subtract the two vectors

jimthompson5910 (jim_thompson5910):

to subtract the two vectors v = (a,b) and w = (c,d), you just follow this rule v - w = (a,b) - (c,d) v - w = (a- c, b - d)

jimthompson5910 (jim_thompson5910):

make sense?

OpenStudy (anonymous):

yes im getting(-3,1)?

jimthompson5910 (jim_thompson5910):

but remember, you're not just subtracting v - w you're really doing 3v - w

jimthompson5910 (jim_thompson5910):

so you're really doing (6, -9) - (5,-2)

jimthompson5910 (jim_thompson5910):

I wrote out v-w as a general template more or less